point charges q1=50uC and q2=-25uC are placed 1m apart. what is the force on a third charge q3 = 20uC play midway between q1 and q2?


1
Expert's answer
2021-05-30T13:29:15-0400
"F_3=k\frac{q_1q_3}{r^2}+k\frac{q_2q_3}{r^2}\\
F_3=9\cdot10^{9}\frac{(50\cdot10^{-6})(20\cdot10^{-6})}{0.5^2}+9\cdot10^{9}\frac{(50\cdot10^{-6})(20\cdot10^{-6})}{0.5^2}"

"F_3=72\ N"